OpenStudy (anonymous):

can someone help me solve this equation ???? 1&1/4 =a+ 3/8

OpenStudy (blurbendy):

1&1/4 = 5/4 5/4 = a + 3/8 a = 5/4 - 3/8 What is 5/4 - 3/8 ?

OpenStudy (anonymous):

7/8

OpenStudy (dumbcow):

correct